### v1

@@ Abstract
Magnetic field inhomogeneity correction remains a challenging task in MRI analysis. Most established techniques are designed for brain MRI by supposing that image intensities in the identical tissue follow a uniform distribution. Such an assumption cannot be easily applied to other organs, especially those that are small in size and heterogeneous in texture (large variations in intensity), such as the prostate. To address this problem, this paper proposes a probabilistic Hadamard U-Net (PHU-Net) for prostate MRI bias field correction. First, a novel Hadamard U-Net (HU-Net) is introduced to extract the low-frequency scalar field, multiplied by the original input to obtain the prototypical corrected image. HU-Net converts the input image from the time domain into the frequency domain via Hadamard transform. In the frequency domain, high-frequency components are eliminated using the trainable filter (scaling layer), hard-thresholding layer, and sparsity penalty. Next, a conditional variational autoencoder is used to encode possible bias field-corrected variants into a low-dimensional latent space. Random samples drawn from latent space are then incorporated with a prototypical corrected image to generate multiple plausible images. Experimental results demonstrate the effectiveness of PHU-Net in correcting bias-field in prostate MRI with a fast inference speed. It has also been shown that prostate MRI segmentation accuracy improves with the high-quality corrected images from PHU-Net. The code will be available in the final version of this manuscript.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI) is an essential tool for clinicians to achieve accurate diagnoses and optimize treatment planning, and several other prognostic tasks. However, the main magnetic field in an MRI scanner is not perfectly uniform across the imaging volume, making the quantification tasks (detection, diagnosis, segmentation) challenging and sub-optimal. To address this non-uniformity problem in MRI, one critical step is to correct bias field  [25, 8]. The bias field refers to a low-frequency multiplicative field arising from various sources, such as coil sensitivity variations and radio-frequency inhomogeneities. It distorts intensity distributions and impacts the accuracy of subsequent downstream tasks, most importantly image segmentation algorithms.
In the past years, numerous approaches have been proposed for bias field removal. The N4 bias field correction algorithm (N4ITK) has gained popularity for its effectiveness in rectifying inhomogeneities within brain MRI by assuming uniform intensity across distinct types of brain tissues [22, 23, 3]. However, this condition is not valid for prostate MRI because there exist large intensity variations within the same prostate tissue. Additionally, N4ITK corrects the bias field with a high computation cost, leading to a low inference speed. To improve inference speed, convolutional layers-based frameworks [19, 21] were proposed. Nevertheless, these methods are limited by their accuracy, efficiency, and generalization ability. Alternatively, adversarial learning [3, 6] has recently emerged as a promising avenue for addressing non-uniformities in medical imaging. This technique contains two models: while a generator corrects bias fields, a discriminator distinguishes corrected images from ground truth images. Although adversarial networks enhance correction accuracy, they still face challenges such as model collapse, sensitivity to hyperparameters, and longer training times.
@@
In this paper, we propose a probabilistic Hadamard U-Net (PHU-Net) to correct the bias field of MRI and we demonstrate its effectiveness in prostate MRI due to its unique challenges of large intensity variations and complexity of tissue interactions. More specifically, our contributions are summarized as follows: (1) We develop a novel Hadamard U-Net based on the trainable filters and hard-thresholding layers to obtain an efficient representation of the underlying MR scalar field. (2) We incorporate a conditional variational autoencoder (CVAE) to model the joint distribution of all intensity values in ground truth and generate extensive plausible bias field corrected images. (3) We design a hybrid loss optimization composed of Kullback Leibler divergence (KLD) loss [16], total variation loss, and mean squared error (MSE) loss to promote sparsity in the learned representations, resulting in improved robustness and generalization to diverse MRI datasets. (4) Through comprehensive experiments on several benchmark datasets, we validate the superior efficacy of PHU-Net compared to state-of-the-art bias field correction techniques, as well as improved segmentation performance. This demonstrates the potential of PHU-Net in enhancing the quality and reliability of medical image analysis pipelines.
@@ 2 Methods
Figure 1: Probabilistic Hadamard U-Net.
The overall framework of the probabilistic Hadamard U-Net (PHU-Net) is presented in Fig. 1. It is a supervised learning network developed for prostate MRI pre-processing. Our pipeline contains two inter-connected modules: a Hadamard U-Net (HU-Net) and a conditional variational autoencoder (CVAE).
@@ 2.1 Hadamard U-Net
where $\mathbf{m}$ and $\mathbf{n}$ are input vectors. $\mathcal{T}(\cdot)$ denotes the HT. *$*_{d}$ is the dyadic convolution. Therefore, the scaling layer in the HT domain works as the convolution filter in the time domain. On the other hand, the scaling layer requires significantly less computational overhead than the time-domain convolutional filtering.
@@ 2.2 Conditional Variational Autoencoder
Since the intensities in the prostate MRI data follow an unknown distribution, the conditional variational autoencoder (CVAE) [9] is applied to establish the joint probability of all intensities in the ground truth $\mathbf{Y}$. As shown in Fig. 1, the prior network maps the raw image $\mathbf{X}$ to a Gaussian distribution parameterized by its mean $\mathbf{u}\in\mathbb{R}^{D}$ and covariance $\mathbf{V}\in\mathbb{R}^{D\times D}$ in the latent space, where $D$ is the dimension of the latent space. Next, the raw image and the ground truth are concatenated together as the input of the posterior network. This posterior network predicts the distribution of bias field-corrected images conditioned on a raw image. In the training step, the random samples $\mathbf{r}_{i}\in\mathbb{R}^{D}$ are drawn from the posterior distribution $F$:
@@
where $p$ stands for the prior network weights. Then, the random samples are incorporated with HU-Net output. Next, the feature fusion block filters all information to obtain the corrected MRI data. Furthermore, a Kullback Leibler divergence (KLD) loss function imposes a penalty on differences between the outputs of the prior and posterior networks.
@@ 3.1.1 Datasets
We use four distinct publicly available MRI prostate datasets (all T2-weighted) in our experiments: HK dataset [13], UCL dataset [13], HCRUDB dataset [11] and AWS dataset [1]. Since N4ITK is one of the most prevalent MRI bias field correction methods in the last decades, we employ it to correct all scans in these datasets and utilize processed scans as a reference baseline. To validate the generalization performance of different methods, all models are trained on the UCL dataset and then tested on the HK, HCRUDB, and AWS datasets. In summary, the training set contains 318 scans, while the three testing sets contain 288, 1,216, and 904 scans, respectively. All these scans are reshaped into $256\times 256$ in-plane resolution while they have different sizes of slices.
@@ 3.1.2 Implementation details
The HU-Net is developed based on 4 convolutional layers. The kernel sizes are $16\times 16$, $7\times 7$, $7\times 7$ and $16\times 16$, respectively. In CVAE, the prior and the posterior nets share a similar encoder structure. This encoder structure is comprised of two $3\times 3$ convolutional blocks with 32 and 64 filters. Each convolutional block consists of 4 convolutional layers activated by the ReLU function. After the first block, a 2$\times$2 average pooling layer with stride 2 downsamples the image tensors. Moreover, the feature fusion block shown in Fig. 1 contains three 1$\times$1 convolutional layers. Each of the first 2 convolutional layers contains 32 filters, and they are activated by the ReLU function. The third convolutional layer contains 1 filter.
Our model is trained using the AdamW optimizer [14] with a batch size of 128 and a learning rate of 0.0001 under 100 training epochs. $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ in Eq. (9) are set to 10, 0.1, 1, and 1, respectively.
@@ 3.1.3 Performance metrics
Metrics
CVnormal-$\downarrow$
SNRnormal-$\uparrow$
Dataset
HK
AWS
HCRUDB
HK
Original
78.24
59.13
53.78
24.64
N4ITK
78.24
56.76
52.41
24.64
CAE
66.97
62.51
51.39
13.18
ITCNN
78.25
59.02
53.69
23.50
PHU-Net
65.95
54.14
43.39
27.05
Figure 2: Comparison between the original and the corrected MRIs on the AWS dataset. Last column is our proposed method.
@@ 3.2 Experimental Results
Table 1 compares the CV results of our proposed PHU-Net versus other state-of-the-art methods on three test datasets. Compared with N4ITK, PHU-Net reduces CV from 78.24 to 65.95 (15.71%), from 56.76 to 54.14 (4.62%), and from 52.41 to 43.39 (17.21%) on the HK, AWS and HCRUDB datasets respectively. PHU-Net also outperforms CAE and ITCNN. These superior performances indicate that PHU-Net has a better capability for removing intensity variations across the images. Additionally, Table 1 also presents SNR values for the MRI dataset corrected by different methods. In this table, models are only evaluated on the HK dataset because other datasets do not contain a clear background as the HK dataset. Background pixel values are essential in SNR calculation. As a result, PHU-Net achieves the highest SNR. Furthermore, only PHU-Net improves SNR compared to the original images, which demonstrates our proposed method introduces no additional noise to the region of interest during the bias field correction stage.
In consequence, experimental results are visualized in Fig. 2. Images processed by PHU-Net exhibit superior intensity uniformity compared to those processed by other baseline methods. This suggests that our proposed method produces the most homogeneous results.
@@
Metrics
CVnormal-$\downarrow$
SNRnormal-$\uparrow$
Dataset
HK
AWS
HCRUDB
HK
No scaling
67.38
54.92
43.12
25.69
No sparsity
71.36
58.60
49.19
25.76
No threshold
67.44
54.83
43.13
25.43
No TV loss
67.21
55.07
43.77
25.73
No CVAE
68.27
56.15
44.09
26.25
PHU-Net
65.95
54.14
43.39
27.05
@@ 3.2.1 Ablation study
To verify each module in PHU-Net, ablation studies are carried out in Table 2. When the scaling layer, sparsity, hard-thresholding layer, TV loss, or CVAE is removed, SNR is reduced by 2.96% at least and 5.99% at most. Although CV improves slightly in the HCRUDB dataset when the scaling layer is not present, it degrades significantly in the HK and AWS datasets. When there is no CVAE, PHU-Net only contains HU-Net. Its results are still better than those of other models as shown in Table 1. It indicates that HU-Net can extract an efficient MR scalar field.
@@
Metrics
Dice(%)
IoU(%)
PPV(%)
Original
71.43
66.22
76.22
N4ITK
72.38
67.38
75.74
CAE
68.20
63.04
74.07
ITCNN
75.51
70.27
80.06
PHU-Net
76.86
71.72
82.34
@@ 3.2.2 Segmentation experiment
Corrected images are further utilized in the segmentation experiments. Here, a traditional probabilistic U-Net [9] is used as the segmentation model. This segmentation model contains a standard U-Net and a CVAE. The top 50% of the corrected dataset is utilized as the training dataset. The rest of the corrected dataset is employed as the testing dataset. The MSE loss function is applied. Other implementation details are the same as in [9]. As shown in Table 3, it achieves the highest Dice, IoU, and PPV on the images corrected by PHU-Net. Therefore, PHU-Net ensures more accurate and reliable segmentation of the underlying tissue structures than other correction models.
@@ 3.2.3 Runtime analyses
Another benefit of PHU-Net lies in its inference efficiency. For one single NIfTI image file which includes 15 slices in the AWS dataset, the bias field correction runtime of N4ITK, CAE, ITCNN, and PHU-Net are around 44.26, 1.06, 1.23, and 1.13 seconds respectively. These tests are executed on a computer with an Intel Core i7-12700H CPU. Therefore, compared to N4ITK, our proposed model exhibits nearly 40 times faster processing speed, and similar inference times with ITCNN and CAE while better in performance than them.
@@ 4 Conclusion
In this paper, we developed a novel framework named PHU-Net for prostate MRI bias field correction. PHU-Net can be divided into two parts: HU-Net and CVAE. The key idea of PHU-Net is to combine a U-Net structure with transform domain methods, trainable filters, and hard-thresholding layers to extract biased field information. Additionally, the random samples drawn by CVAE were incorporated with PHU-Net results to generate plausible corrected images. Experimental results showed that PHU-Net achieved a better bias field correction performance than other state-of-the-art methods with a fast running speed. What is more, PHU-Net significantly improved the segmentation accuracy of the probabilistic U-Net.
{credits}
4.0.1 Acknowledgements
*** was supported by ***

### v2

@@ Abstract
Magnetic field inhomogeneity correction remains a challenging task in MRI analysis. Most established techniques are designed for brain MRI by supposing that image intensities in the identical tissue follow a uniform distribution. Such an assumption cannot be easily applied to other organs, especially those that are small in size and heterogeneous in texture (large variations in intensity) such as the prostate. To address this problem, this paper proposes a probabilistic Hadamard U-Net (PHU-Net) for prostate MRI bias field correction. First, a novel Hadamard U-Net (HU-Net) is introduced to extract the low-frequency scalar field. HU-Net converts the input image from the time domain into the frequency domain via Hadamard transform. In the frequency domain, high-frequency components are eliminated using the trainable filter (scaling layer), hard-thresholding layer, and sparsity penalty. Next, a conditional variational autoencoder is used to encode possible bias field-corrected variants into a low-dimensional latent space. Random samples drawn from latent space are then incorporated with a prototypical corrected image to generate multiple plausible images. Experimental results demonstrate the effectiveness of PHU-Net in correcting bias-field in prostate MRI with a fast inference speed. It has also been shown that prostate MRI segmentation accuracy improves with the high-quality corrected images from PHU-Net. The codes are available at https://github.com/Holmes696/Probabilistic-Hadamard-U-Net.
 [^1] Corresponding auhtor’s email: . This work is supported by the NIH funding: R01-CA246704, R01-CA240639, U01-DK127384-02S1, and U01-CA268808, and NSF IDEAL 2217023.
@@ 1 Introduction
The main magnetic field in an Magnetic Resonance Imaging (MRI) scanner is not perfectly uniform across the imaging volume, making the quantification tasks (detection, diagnosis, segmentation) challenging and sub-optimal. To address this non-uniformity problem in MRI, one critical step is to correct bias field  [25, 8]. The bias field refers to a low-frequency multiplicative field arising from various sources, such as coil sensitivity variations and radio-frequency inhomogeneities. It distorts intensity distributions and impacts the accuracy of subsequent downstream tasks, most importantly image segmentation algorithms.
Numerous approaches have been proposed for bias field removal. Among them, the N4 bias field correction algorithm (N4ITK) has gained popularity for its effectiveness in rectifying inhomogeneities within brain MRI by assuming uniform intensity across distinct types of brain tissues [22, 23, 3]. However, this condition is not valid for prostate MRI because there exist large intensity variations within the same prostate tissue. Additionally, N4ITK corrects the bias field with a high computation cost, leading to a low inference speed. To improve inference speed, convolutional layers-based frameworks [19, 21] were proposed. Yet, these methods are limited by their accuracy, efficiency, and generalization ability. Alternatively, adversarial learning [3, 6] has emerged as a promising avenue for addressing non-uniformities in medical imaging. This technique contains two models: while a generator corrects bias fields, a discriminator distinguishes corrected images from ground truth images. Although adversarial networks enhance correction accuracy, they still face challenges such as model collapse, sensitivity to hyperparameters, and longer training times.
@@
In this study, we propose a probabilistic Hadamard U-Net (PHU-Net) to correct the bias field of MRI and we demonstrate its effectiveness in prostate MRI due to its unique challenges of large intensity variations and complexity of tissue interactions. More specifically, our contributions are summarized as follows: (1) We develop a novel Hadamard U-Net based on the trainable filters and hard-thresholding layers to obtain an efficient representation of the underlying MR scalar field. (2) We incorporate a conditional variational autoencoder (CVAE) to model the joint distribution of all intensity values in ground truth and generate extensive plausible bias field corrected images. (3) We design a hybrid loss optimization composed of Kullback Leibler divergence (KLD) loss [16], total variation loss, and mean squared error (MSE) loss to promote sparsity in the learned representations, resulting in improved robustness and generalization to diverse MRI datasets. (4) Through comprehensive experiments on several benchmark datasets, we validate the superior efficacy of PHU-Net compared to state-of-the-art bias field correction techniques, as well as improved segmentation performance. This demonstrates the potential of PHU-Net in enhancing the quality and reliability of medical image analysis pipelines.
Figure 1: Probabilistic Hadamard U-Net.
@@ 2 Methods
The overall framework of the PHU-Net is presented in Fig. 1. Our supervised pipeline contains two inter-connected modules: a Hadamard U-Net (HU-Net) and a conditional variational autoencoder (CVAE).
@@ 2.1 Hadamard U-Net
where $\mathbf{m}$ and $\mathbf{n}$ are input vectors. $\mathcal{T}(\cdot)$ denotes the HT. $*_{d}$ is the dyadic convolution. Therefore, the scaling layer in the HT domain works as the convolution filter in the time-domain. On the other hand, the scaling layer requires significantly less computational overhead than the time-domain convolutional filtering.
@@ 2.2 Conditional Variational Autoencoder
Since the intensities in the prostate MRI data follow a non-uniform distribution, the conditional variational autoencoder (CVAE) [9] is applied to establish the joint probability of all intensities in the ground truth $\mathbf{Y}$. As shown in Fig. 1, the prior network maps the raw image $\mathbf{X}$ to a Gaussian distribution parameterized by its mean $\mathbf{u}\in\mathbb{R}^{D}$ and covariance $\mathbf{V}\in\mathbb{R}^{D\times D}$ in the latent space, where $D$ is the dimension of the latent space. Next, the raw image and the ground truth are concatenated together as the input of the posterior network. This posterior network predicts the distribution of bias field-corrected images conditioned on a raw image. In the training step, the random samples $\mathbf{r}_{i}\in\mathbb{R}^{D}$ are drawn from the posterior distribution $F$:
@@
where $p$ stands for the prior network weights. Then, the random samples are incorporated with HU-Net output. Next, the feature fusion block filters all information to obtain the corrected MRI data. Furthermore, a KLD loss function imposes a penalty on differences between the outputs of the prior and posterior networks.
@@ 3.1.1 Datasets
We use four distinct publicly available MRI prostate datasets (all T2-weighted) in our experiments: HK dataset [13], UCL dataset [13], HCRUDB dataset [11] and AWS dataset [1]. Since N4ITK is one of the most prevalent MRI bias field correction methods in the last decades, we employ it to clean all scans in these datasets and utilize processed scans as a reference baseline that serves as a substitute for ground truth. To validate the generalization performance of different methods, all models are trained on the UCL dataset and then tested on the HK, HCRUDB, and AWS datasets. In summary, the training set contains 318 scans, while the three testing sets contain 288, 1,216, and 904 scans, respectively. All these scans are reshaped into $256\times 256$ in-plane resolution while they have different sizes of slices.
@@ 3.1.2 Implementation details
The HU-Net includes 4 convolutional layers. The kernel sizes are $16\times 16$, $7\times 7$, $7\times 7$ and $16\times 16$, respectively. In CVAE, the prior and the posterior nets share a similar encoder structure. This encoder structure is comprised of two $3\times 3$ convolutional blocks with 32 and 64 filters. Each convolutional block consists of 4 convolutional layers activated by the ReLU function. After the first block, a 2$\times$2 average pooling layer with stride 2 downsamples the image tensors. Moreover, the feature fusion block shown in Fig. 1 contains three 1$\times$1 convolutional layers. Each of the first 2 convolutional layers contains 32 filters, and they are activated by the ReLU function. The third convolutional layer contains 1 filter. Our model is trained using the AdamW optimizer [14] with a batch size of 128 and a learning rate of 0.0001 under 100 training epochs. $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ in Eq. (9) are set to 10, 0.1, 1, and 1, respectively.
@@ 3.1.3 Performance metrics
[…]
Dataset HK AWS HCRUDB HK
Original 78.24 59.13 53.78 24.64
N4ITK 78.24 56.76 52.41 24.64
CAE 66.97 62.51 51.39 13.18
ITCNN 78.25 59.02 53.69 23.50
PHU-Net 65.95 54.14 43.39 27.05
Figure 2: Comparison between the original and the corrected MRIs on the AWS and HCRUDB dataset. Last column is our proposed method.
@@ 3.2 Experimental Results
Table 1 compares the CV results of our proposed PHU-Net versus other state-of-the-art methods on three test datasets. Compared with N4ITK, PHU-Net reduces CV from 78.24 to 65.95 (15.71%), from 56.76 to 54.14 (4.62%), and from 52.41 to 43.39 (17.21%) on the HK, AWS and HCRUDB datasets, respectively. PHU-Net also outperforms CAE and ITCNN, indicating that PHU-Net has a better capability for removing intensity variations regardless of scanner types. Additionally, Table 1 also presents SNR values for the MRI dataset corrected by different methods. In this comparison, models are only evaluated on the HK dataset because other datasets do not contain a clear background as the HK dataset, essential for SNR calculation. As a result, PHU-Net achieves the highest SNR. Furthermore, only PHU-Net improves SNR compared to the original images, which demonstrates our proposed method introduces no additional noise to the region of interest during the bias field correction stage.
Experimental results are visualized in Fig. 2. Images processed by PHU-Net exhibit superior intensity uniformity compared to those processed by other baseline methods.
@@
[…]
Dataset HK AWS HCRUDB HK
No scaling 67.38 54.92 43.12 25.69
No sparsity 71.36 58.60 49.19 25.76
No threshold 67.44 54.83 43.13 25.43
No TV loss 67.21 55.07 43.77 25.73
No CVAE 68.27 56.15 44.09 26.25
PHU-Net 65.95 54.14 43.39 27.05
@@ 3.2.1 Ablation study
To verify each module in PHU-Net, ablation studies are carried out in Table 2. When the scaling layer, sparsity, hard-thresholding layer, TV loss, or CVAE are removed, SNR is reduced by 2.96% at least and 5.99% at most. Although CV improves slightly in the HCRUDB dataset when the scaling layer is not present, it degrades significantly in the HK and AWS datasets. When there is no CVAE, PHU-Net only contains HU-Net. Its results are still better than those of other models in Table 1.
@@
Metrics Dice(%) IoU(%) PPV(%)
Original 71.43 66.22 76.22
N4ITK 72.38 67.38 75.74
CAE 68.20 63.04 74.07
ITCNN 75.51 70.27 80.06
PHU-Net 76.86 71.72 82.34
@@ 3.2.2 Segmentation experiment
Corrected images are further utilized in the segmentation experiments. Here, a traditional probabilistic U-Net [9] is used as the segmentation model, containing a standard U-Net and a CVAE. The top 50% of the corrected dataset is utilized as the training dataset while the remaining is employed as the testing dataset. The MSE loss function is applied. Other implementation details are the same as in [9]. As shown in Table 3, it achieves the highest Dice, IoU, and PPV on the images corrected by PHU-Net. Therefore, PHU-Net ensures more accurate and reliable segmentation of the underlying tissue structures than other correction models.
@@ 3.2.3 Runtime analyses
Another benefit of PHU-Net lies in its inference efficiency. For one single NIfTI image file ( 1̃5 slices), the bias field correction runtime of N4ITK, CAE, ITCNN, and PHU-Net are around 44.26, 1.06, 1.23, and 1.13 seconds, respectively. These tests are executed on a computer with an Intel Core i7-12700H CPU. Compared to N4ITK, our proposed model exhibits nearly 40 times faster processing speed, and similar inference times with ITCNN and CAE while better in performance than them.
@@ 4 Conclusion
In this paper, we developed a novel framework named PHU-Net for MRI bias field correction. The key idea of PHU-Net is to combine a U-Net structure with transform domain methods, trainable filters, and hard-thresholding layers to extract biased field information. Experimental results showed that PHU-Net achieved a better bias field correction performance than other state-of-the-art methods with a fast running speed. What is more, PHU-Net significantly improved the segmentation accuracy of the probabilistic U-Net.